OpenStudy (anonymous):

HELP ME FIND THE ZEROS I got super lost and confused myself y=3sin(x-(pi/4))

OpenStudy (anonymous):

simplify?

OpenStudy (jdoe0001):

recall from the transformations \(\large \begin{array}{llll} y=3sin&\left(x-\frac{\pi}{4}\right)\\ &\qquad \uparrow \\ &\textit{horizontal shift} \end{array}\)

OpenStudy (jdoe0001):

so the zeros will be the same as for sin(x), just shifted over to the right that much

OpenStudy (anonymous):

soo how do I write that in an equation I think it goes something like +/- pi/4 n

OpenStudy (jdoe0001):

hmmm well, the zeros will be when the graph touches the y-axis so you're just being asked on what is the value of "x" at those points, that is, what angles find them on the parent function of sin(x) and shift them horizontally THAT MUCH to get it for this function

OpenStudy (anonymous):

ahhhhh im so sorry you are giving sucha great explanation I am just soo lost

OpenStudy (jdoe0001):

graph sin(x) notice where the zeros happen then shift that by \(\bf \frac{\pi}{4}\) to the left

OpenStudy (jdoe0001):

ahemm to the right rather

OpenStudy (anonymous):

is the answer 1/3+ pi/4 N

OpenStudy (anonymous):

@jdoe0001